# Overestimation of melting temperatures calculated by first-principles molecular dynamics simulations

### v1

@@ III.2 Silicon
Figure 3: The diffusion coefficient, $D^{\prime}$, and structural energy, $E_{\rm st}$, of Na calculated by NC+GGA. Cell sizes are $N=16$ and 128. The $k$ meshes are $2^{3}$ for $N=16$, whereas are $1$ (red circles) and $2^{3}$ (blue circles) for $N=128$. $E_{\rm cut}=245$ eV. The experimental value for $T_{m}$ is indicated by the dashed line. For each $N$, the energy origin was set to the ground state energy for that cell.
@@
Cell sizes of 16 and 128 atoms were examined. In Fig. 3, the structural energy, $E_{\rm st}$, and diffusion constant, $D^{\prime}$, are plotted as functions of $T$. Let us first check the cell size dependence on $T_{m}$. When different sizes of supercells are compared, the equivalent $k$ meshes should be use. Thus, the data obtained by a $2^{3}$ $k$-mesh of 16-atom cell (crosses in the figure) should be compared with the data obtained by a $1$ $k$-mesh of 128-atom (red circles). Clearly, the $T_{m}$ value is lowered from 600 to 460 K by increasing the cell size from 16 to 128. Accordingly, we can expect further decrease in the $T_{m}$ value, approaching the experimental value 371 K, by increasing the cell size.
For metallic systems, the $k$ mesh can affect the total energy calculation largely. Hence, the dependence of the $k$ mesh was also examined for the 128-atom cell. However, there was only a marginal change in $T_{m}$ between $1$ (red circles) and $2^{3}$ (blue circles) meshes, as shown in this figure.
Although a non-negligible overlap between the core and valence electrons can require a correction in some cases Louie et al. (1982), there is no difficulty in constructing the pseudopotential for this case. Therefore, the primary issue causing the incorrect $T_{m}$ for Na appears to be the slow convergence against the cell size.
III.4 Silica
@@
The last example is an oxide material, namely, crystalline $\alpha$-quartz. The authors previously used this crystal as the parent material for the formation of silica glass and found the overestimation of $T_{m}$ Shirai et al. (2023). The primitive unit cell of $\alpha$-quartz has the rhombohedral Bravais lattice comprising three SiO2 units. The data for $\alpha$-quartz listed in Table 1 represent the results of MD simulations performed by the Phase code. These calculations employed the US+GGA with 72-atom cell having dimensions of $2\times 2\times 2$. Unfortunately, both codes Phase and VASP failed to keep long time MD simulations with a smaller cell having dimension of $1\times 1\times 1$. Under these conditions, a drastic error propagations in atom forces occurred at some steps, leading to a break in the simulation. Osaka2k code using NC-LDA was found to be stable against the long run of MD simulations using this cell size.not (c) Hence, the cell-size dependence for $\alpha$-quartz was examined by osaka2k code with the NC-LDA potential, using $E_{\rm cut}$=544 eV for the planewave expansion. The results are shown in Fig. 4. Although this cutoff energy is insufficient to achieve convergence of the total energy for oxide materials, the value represents a compromise required to allow MD simulations with the NC potentials. As a consequence of this low cutoff energy, the potential of oxygen becomes effectively softer, leading to lowering $T_{m}$. This effect led to the reduced value of 3500 K in Fig. 4 compared with that of 4500 K in Table 1. Even so, the value of $T_{m}=3500$ K is still much higher than the experimental result of 1700 K.
@@
Figure 4: The diffusion coefficient, $D^{\prime}$, and structural energy, $E_{\rm st}$, for $\alpha$-quartz calculated by NC+LDA. $E_{\rm cut}=544$ eV. p1 indicates results obtained with a $1\times 1\times 1$ cell ($4^{3}$ $k$-mesh), while p8 indicates a $2\times 2\times 2$ cell ($1$). p8’ indicates the same cell with a volume expansion of 16.9 %. The experimental value of $T_{m}$ is indicated by the dashed line. For each $N$, the energy origin was set to the ground state energy of that cell. For p8’, it was set to that for the p8 cell.
@@
As shown in Fig. 4, the onset of melting almost does not change upon increasing the cell size from 9 to 72 atoms. The calculated $T_{m}$ seems rather to be saturated at the large overestimated value. Similar results were obtained for $\beta$-cristobalite in that increasing the unit cell dimension from $1\times 1\times 1$ to $2\times 2\times 2$ did not alter the overestimated $T_{m}$ (see Supplemental Materials). Furthermore, the effect of volume expansion was also examined. The actual volume of the liquid silica is larger than that of the crystalline $\alpha$-quartz by 16.9 %. The same $2\times 2\times 2$ cell was assessed in conjunction with this volume expansion, referred to as the p8’ data in Fig. 4. Even with this large volume expansion, there was no discernible change in $T_{m}$. Although the possibility of reducing $T_{m}$ by further increasing the cell size cannot be excluded, it is unlikely that the large overestimation ratio, $r$, would be fixed by this change.
@@
We rather consider that the overestimation of $T_{m}$ for $\alpha$-quartz is primarily due to the error of the potential. The systematic error of LDA is well known in that LDA overestimates the binding energy and accordingly underestimates the lattice parameters about a few % Jones and Gunnarsson (1989). The LDA error is a result of adapting the energy functional obtained for homogeneous electron systems to localized systems. Hence, the error generally becomes larger for strongly localized states. The underestimation of lattice parameters in LDA is fixed by GGA Jones and Gunnarsson (1989); Perdew et al. (1996); Zhang and Yang (1998); Perdew and Kurth (2003). However, the overestimate of the binding energy still remains in GGA, although the error is largely improved. Oxygen atom has the largest error. Patton et al. report that the overbinding $\Delta$ by GGA is 0.53 eV for O2 molecule Patton et al. (1997). For H2O, $\Delta$ is only 0.03 eV, but this amounts to a change in $T_{m}$ by 300 K. When calculating the formation energies of metal oxides, the formation energy for the oxygen part is typically adjusted by approximately 0.7 eV/(O atom) Wang et al. (2006); Jain et al. (2011); Mutter et al. (2020).
@@
The cohesive energy, $E_{\rm coh}$, is the energy difference between an isolated atom and the solid state, representing the extrema of localized and extended states, respectively. Therefore, the LDA/GGA error appears in $E_{\rm coh}$ most significantly. Because $H_{m}$ is the energy difference between the solid and liquid states, a similar effect is expected. Here is, however, a subtle problem of the electron correlation in liquids. Normally, the electron system having a high atom density represents the extended state (solid state), while an electron system having a low atom density represents the localized state (atom state). However, the average density of a liquid is close to that of the solid for a given material: the difference is only a few %. There are even cases, such as Si, in which the liquid state is denser than the solid state. Therefore, the average density and average atom distance are not suitable to estimate the degree of localization of electrons for the case of liquid. Rather, in Secs. II.1 and III.2, we have seen that energy dissipation processes play important roles in describing the thermodynamic properties of liquids.
@@
The energy dissipation can be described by the time correlation in a given pair of atoms (Egelstaff (1992), Chap. 9),
@@
The bar indicates the time average with respect to $t^{\prime}$. The distance between atoms in a particular atom pair, $R_{ij}$, immediately increases with time after melting. This correlation function may be approximately presented by $G_{ij}({\bf r},t)\approx e^{-t/\tau_{ij}}\delta({\bf r}-\bar{d})$, where $\bar{d}$ is the average atom distance. The first factor $e^{-t/\tau_{ij}}$ is important. The relaxation time $\tau_{ij}$ is determined by the energy barrier between atoms $i$ and $j$. This corresponds to the saddle point of the energy landscape. In contrast, RDF is obtained differently. The correlation is evaluated first by taking the particle average and then by taking the time average, as
@@
This correlation, and accordingly, RDF reflect only the average atom distance, which corresponds to the minima in the energy landscape. The information of energy dissipation is completely missed, and thus the diffusing nature of liquid is masked in RDF. Agreement not only in the energy minima but also in the saddle points is needed to reproduce the thermodynamic properties of liquids. Consequently, even though the averaged density does not change significantly on melting, it is highly likely that a significant change occurs in the electron correlation at $T_{m}$. Thus, the LDA/GGA error also affects the energy barrier between the solid and liquid states, leading to an overestimation of $T_{m}$.
@@
Lastly, it is worthy to mention a distinct feature of the liquid phase of silica. Normally, significant changes in $E_{\rm st}$ occur only at phase transitions: in this study, near $T_{m}$. However, as shown in Fig. 4, for the silica liquid, an apparent increase in $E_{\rm st}$ with increasing $T$ persists over the entire $T$ range examined. This implies that the silica liquid is comprised of numerous distinct metastable states, which are separated each other by energy barriers. This could explain why the silica liquid has the highest viscosity among various liquids. However, many things about the thermodynamic properties of silica liquid still remain unknown.
IV Conclusion
This work demonstrated systematic overestimations of the melting temperature in FP-MD simulations. Two important factors, the size dependence of the cell and the LDA/GGA error, were studied. The importance of these factors vary depending on the material under consideration. The convergence with respect to the cell size is not fast, unless the potential is very shallow. The significant finding of this study is demonstration of the effect of the LDA/GGA error on $H_{m}$ and $T_{m}$. This error cannot be removed by increasing the cell size.
Throughout this study, correct understanding thermodynamic nature of liquids is emphasized. There is no eigenstate for liquids, and instead the thermodynamic equilibrium of a liquid is established by the balance between incessant excitation and deexcitation in atom motions. The phonon model often hides this important nature of liquids. The total energy in adiabatic MD simulations describes all the process of energy relaxation, and hence gives the correct relationship between $U$ and $T$.
Acknowledgment

### v2

@@ 3.3 Sodium
Figure 3: The diffusion coefficient, $D$, and structural energy, $E_{\rm st}$, of Na calculated by NC+GGA. Cell sizes are $N=16$ and 128. The $k$ meshes are $2^{3}$ for $N=16$, whereas are $1$ (red circles) and $2^{3}$ (blue circles and black diamonds) for $N=128$. All the data except marked by diamond were obtained using the cell volume of crystal ($V_{\rm sol}$). Diamond data indicate using the liquid volume ($V_{\rm liq}$). $E_{\rm cut}=245$ eV. The experimental value for $T_{m}$ is indicated by the dashed line. For each $N$, the energy origin was set to the ground state energy for that cell.
@@
Cell sizes of 16 and 128 atoms were examined. In Fig. 3, the structural energy, $E_{\rm st}$, and diffusion constant, $D$, are plotted as functions of $T$. In this case, even though a finite width of the transition region is observed, there is no region of negative specific heat. Let us first check the cell size dependence on $T_{m}$. When different sizes of supercells are compared, the equivalent $k$ meshes should be employed. Thus, the data obtained by a $2^{3}$ $k$-mesh of 16-atom cell (crosses in the figure) should be compared with the data obtained by a $1$ $k$-mesh of 128-atom (red circles). Clearly, the $T_{m}$ value is lowered from 600 to 460 K by increasing the cell size from 16 to 128. Accordingly, we can expect further decrease in the $T_{m}$ value, approaching the experimental value 371 K, by increasing the cell size.
For metallic systems, the $k$ mesh can affect the total energy calculation largely. Hence, the dependence of the $k$ mesh was also examined for the 128-atom cell. However, there was only a marginal change in $T_{m}$ between $1$ (red circles) and $2^{3}$ (blue circles) meshes, as shown in this figure. Finally, the effect of cell volume was examined. The volume of the liquid phase is expanded from that of the solid by 9.4 %. The date marked by diamonds in the figure indicate the results of the melting simulations on the expanded unit cells. The effect of volume expansion is apparent. If we determine $T_{m}$ by the onset of the diffusion coefficient, we already obtain a good agreement with the experimental value, $T_{m}=371$ K. The agreement in $T_{m}$ was also reported by Raty et al. [88] for the pressure dependence of $T_{m}$ of Na: they used $N=128$. They reported that the agreement was obtained without use of the two-phase coexistence method; however, the agreement is not clear at $p=0$. This agreement in $T_{m}$ further confirms that the presence of surface/interface is not a universal requirement for calculation of $T_{m}$ for bulk materials.
There is a view that the equilibrium concentration of lattice vacancies reduces the melting temperature of bulk crystals through an increase in the melting entropy. The overestimation of $T_{m}$ by MD simulation is sometimes ascribed to this effect. This is tantamount to say that the observed $T_{m}$ is not the property of the perfect crystal. From the historical perspective, the effect of vacancies near $T_{m}$ was proposed for sodium by observing the correlation between the increases in the specific heat and thermal expansivity near $T_{m}$ [89, 90]. The formation energy of vacancy was estimated as 0.12-0.16 eV, and hence non-negligible concentrations of vacancies are certainly present at $T_{m}$. However, today, it is recognized that these increases are better described by the anharmonic phonon effects. This vacancy mechanism is often repeated to correct the overestimation of $T_{m}$. For example, Chaplot introduced 1 % vacancies for MgSiO3 to reduce the calculated $T_{m}$ [91]: however, see a criticism by Belonoshko [33]. For MgO, Arkhipin et al. introduced about 20 % of voids to reduce $T_{m}$ by 500 K [92]. Such high concentrations of vacancy are unrealistic, unless the formation energy of defects is very small.
3.4 Silica
@@
The last example is an oxide material, namely, crystalline $\alpha$-quartz. The authors previously used this crystal as the parent material for the formation of silica glass and found the overestimation of $T_{m}$ [28]. The primitive unit cell of $\alpha$-quartz has the rhombohedral Bravais lattice comprising three SiO2 units. The data for $\alpha$-quartz listed in Table 1 represent the results of MD simulations performed by the Phase code. Unfortunately, both codes Phase and VASP failed to keep long time MD simulations with a smaller cell having dimension of $1\times 1\times 1$. Drastic error propagations in atom forces occurred at some steps, leading to a break in the simulation. Osaka2k code using NC-LDA was found to be stable against the long run of MD simulations using this cell size.[93] Hence, the cell-size dependence for $\alpha$-quartz was examined by osaka2k code with the NC-LDA potential, using $E_{\rm cut}$=544 eV for the planewave expansion. The results are shown in Fig. 4. Although this cutoff energy is insufficient to achieve convergence of the total energy for oxide materials, it is meaningful for the purpose of examining the size dependence. As the result of this low cutoff energy, the potential of oxygen becomes effectively softer, which makes it easier to test the size dependence. The calculated value $T_{m}$ is reduced to 3500 K, as shown in Fig. 4, compared with that of 4500 K in Table 1. Even so, the value of $T_{m}=3500$ K is still much higher than the experimental result of 1700 K.
@@
Figure 4: The diffusion coefficient, $D$, and structural energy, $E_{\rm st}$, for $\alpha$-quartz calculated by NC+LDA. $E_{\rm cut}=544$ eV. p1 indicates results obtained with a $1\times 1\times 1$ cell ($4^{3}$ $k$-mesh), while p8 indicates a $2\times 2\times 2$ cell ($1$). p8’ indicates the same cell with a volume expansion of 16.9 %. The experimental value of $T_{m}$ is indicated by the dashed line. For each $N$, the energy origin was set to the ground state energy of that cell. For p8’, it was set to that for the p8 cell.
@@
As shown in Fig. 4, the onset of melting almost does not change upon increasing the cell size from 9 to 72 atoms. The calculated $T_{m}$ seems rather to be saturated at the large overestimated value. Similar results were obtained for $\beta$-cristobalite in that increasing the unit cell dimension from $1\times 1\times 1$ to $2\times 2\times 2$ did not alter the overestimated $T_{m}$ (see Supplemental Materials). Generally, the size dependence is significant for small $N$ and is reduced as $N$ increases: for example, see Figs. 5-7 of model potentials by Hong and van de Walle [17]. The behavior of almost no change in $T_{m}$ between p1 and p8 implies that no significant reduction is likely to occur by further increase of $N$. Furthermore, the effect of volume expansion was examined. The actual volume of the liquid silica is larger than that of the crystalline $\alpha$-quartz by 16.9 %. The expanded cell by this amount was used, whose results are indicated by the p8’ data in Fig. 4. Even with this large volume expansion, there is no discernible change in $T_{m}$. This is a contrasting behavior to the Na case. This insensitivity to the volume expansion suggests that there is some strong cause to fix the overestimation of $T_{m}$, irrespective of the cell volume as well as $N$.
@@
We consider that the primal cause of the overestimation of $T_{m}$ for $\alpha$-quartz is the error of the potential. The systematic error of LDA is well known in that LDA overestimates the binding energy and accordingly underestimates the lattice parameters about a few % [18]. The LDA error is a result of adapting the energy functional obtained for homogeneous electron systems to localized systems. Hence, the error generally becomes larger for strongly localized states. The underestimation of lattice parameters in LDA is fixed by GGA [18, 94, 95, 96]. In fact, in the calculation of $T_{m}$ of MgO by the coexistence method, Alfé showed a reduction of $T_{m}$ by GGA compared to LDA [3] [97]: the amount of the reduction is about 500 K. His result already indicates that the calculated $T_{m}$ value varies depending on the potential even within DFT. The reduction in $T_{m}$ by GGA, however, does not guarantee that GGA is good enough. The overestimate of the binding energy still remains in GGA. Oxygen atom has the largest error. Patton et al. report that the overbinding $\Delta$ by GGA is 0.53 eV for O2 molecule [98]. For H2O, $\Delta$ is only 0.03 eV, but this amounts to a change in $T_{m}$ by 300 K. When calculating the formation energies of metal oxides, it is a common practice to adjust the oxygen contribution by setting approximately 0.7 eV/(O atom) [99, 100, 101]. Presently, there is no way to cope this adjustment in the MD simulation of melting.
@@
The cohesive energy, $E_{\rm coh}$, is the energy difference between an isolated atom and the solid state, representing the extrema of localized and extended states, respectively. Therefore, the LDA/GGA error appears in $E_{\rm coh}$ most significantly. Because $H_{m}$ is the energy difference between the solid and liquid states, a similar effect is expected. Here is, however, a subtle problem of the electron correlation in liquids. Normally, the electron system having a high atom density represents the extended state (solid state), while an electron system having a low atom density represents the localized state (atom state). However, the average density of a liquid is close to that of the corresponding solid: the difference is only a few %. There are even cases, such as Si, in which the liquid state is denser than the solid state. Therefore, the average density and average atom distance are not suitable to estimate the degree of localization of electrons for the case of liquid. Rather, in Sec. 2.1, we emphasized the importance of the time correlation in atom positions.
@@
The energy dissipation can be described by the time correlation in a given pair of atoms ([47], Chap. 9),
@@
The bar indicates the time average with respect to $t^{\prime}$. The distance between those atoms in a particular atom pair, $R_{ij}$, immediately increases with time after melting. This correlation function may be approximately presented by $G_{ij}({\bf r},t)\approx e^{-t/\tau_{ij}}\delta({\bf r}-\bar{d})$, where $\bar{d}$ is the average atom distance. The first factor $e^{-t/\tau_{ij}}$ is important. The relaxation time $\tau_{ij}$ is determined by the energy barrier between atoms $i$ and $j$. This corresponds to the saddle point of the energy landscape. In contrast, RDF is obtained differently. The correlation is evaluated first by taking the particle average and then by taking the time average, as
@@
This correlation reflects only the average atom distance, which corresponds to the minima in the energy landscape. The information of energy dissipation is lost, and thus the diffusing nature of liquid is masked in RDF. Agreement not only in the energy minima but also in the saddle points is needed to reproduce the thermodynamic properties of liquids. Consequently, even though the averaged density does not change significantly on melting, it is likely that a significant change occurs in the electron correlation. Thus, the LDA/GGA error also affects the energy barrier between the solid and liquid states, leading to an overestimation of $T_{m}$.
@@
Lastly, it is worthy to mention a distinct feature of the liquid phase of silica. Normally, significant changes in $E_{\rm st}$ occur only at phase transitions. However, as shown in Fig. 4, for the silica liquid, an apparent increase in $E_{\rm st}$ with increasing $T$ persists over the entire temperature range examined. This implies that the silica liquid is comprised of numerous distinct metastable states, which are separated each other by energy barriers. This could explain why the silica liquid has the highest viscosity among various liquids. Even in the liquid of silica, hysteresis is observed in the specific heat measurement [102]. However, many things about the thermodynamic properties of silica liquid still remain unknown.
4 Conclusion
The difficulty in calculating $T_{m}$ consists in the thermodynamic nature of liquids. The thermodynamic temperature is determined by the time average of the kinetic energy, through Eq. (9). This average is influenced by the atom coordinates at equilibrium. For solids, the equilibrium atom positions are unique and $T$ is determined by the bottom part of potential only. The equilibrium state is insensitive to the relaxation processes in which that state is led. In contrast, for liquids, the atom coordinate to determine the equilibrium continuously changes, and the temperature is determined by transitions between different coordinates. The details of energy relaxation processes thus affects $T_{m}$. The relaxation is controlled by the energy barriers, which are saddle points in the potential. One implication of this is that this explains why two similar potentials that give the almost same thermodynamic properties yield very different values for $T_{m}$, unless the two potentials are constructed so as to match up to the saddle points. Another one is that introduction of artificial heat bath should be avoided, because it alters the native balance among different relaxations. Adiabatic MDs are recommended.
The present simulations based on the adiabatic FP-MD resolved the presence of the transition width $W_{m}$, which is unavoidable for finite-size cells. The width can be removed only when the thermodynamic limit, $N\rightarrow\infty$, is taken. In particular, this width becomes harmful when the transition region exhibits negative specific heat. The extrema in the transition region could produce spurious equilibria. Bearing these cautions in mind, we have deduced the following conclusions from the present simulations. First, a good agreement in $T_{m}$ is obtained for Si and Na by the adiabatic FP-MD simulations up to the uncertainty $W_{m}$. The fact that these agreements are obtained by the modest size of cells, approximately 100 atoms, encourages us to use the standard method of FP-MD simulations. This also implies that the involvement of surface in determining $T_{m}$ for bulk solids cannot be a universal reasoning of use of the two-phase coexistence method, unless the surface energy or other factors are too large. Second, there are many cases in which the adiabatic FP-MD simulations give large overestimations for $T_{m}$. These cases are mainly those materials containing oxygen atoms. It is highly likely that the overestimation in $T_{m}$ is brought about by the LDA/GGA error in the electron correlation functional. The strong insensitivity against the change in $N$ and in the cell volume supports this conclusion.